8x+4=2. Therefore x=?

8x+4=2.
Therefore x=?

Grade:12th Pass

58 Answers

Sunil Kumar FP
9 years ago
we have, 8x+4=2
adding -4 to both side,we get
(8x+4)-4=2-4
8x+4-4=-2
8x=-2
8x/8=-2/8
x=-1/4
thus the value of x=-1/4 for which the equation is satisfied.
check
LHS= 8x+4 =8*(-1/4)+4 =2
RHS=2
Therefore left hand side=right hand side
Nihal Patel
9 years ago
8x+4=2
8x=4-2
x=2/8
Therefor, x=1/3
 
H KRISHNA
9 years ago
8x+4=2
8x=2-4
8x=-2
x=-2/8
:. x=-1/4
 
Mohammed Anas
9 years ago
U take the like terms and then u get.
8x=2-4
8x=-2
Then divide both sides by 8 to get the value of x
Therfore,
x=-2/8
Simplify that:
x=-1/4
 
FINAL ANS: x=-1/4
Aakash
8 years ago
\frac{-1}{4}
noogler
8 years ago
-1/4
suman gupta
8 years ago
-1/4
 
Prajwal Kavad
8 years ago
x=-1/4
mamam
8 years ago
it is differentiation and  it is too high level for 6th grade
VAISHNAVI BATHINA
8 years ago
(8x+4)-4=2-48x+4-4=-28x=-28x/8=-2/8x=-1/4
Adarsh
8 years ago
x= -1/4 is the answer for the given question
Nandini Menon
8 years ago
8x+4=2
8x=2-4  (transposition)
8x= -2
x= -2/8
therefore, x= -1/4 (because -2/8 is divided into the lowest form)
 
Hinal Srivastava
8 years ago
 
8x+4=2
8x=4-2
x=2/8
Therefor, x=1/3
SAHIL
8 years ago
-1/4
OUTPASS
8 years ago
8x+4=2,
8x=2-4,
8x= -2,
x =-2/4
   =-1/4
Therefore the value of x is -1/4
pralay
8 years ago
-1/2
OUTPASS
8 years ago
-1/4
SHANMUKESHWAR
8 years ago
8x+4=2
8x=-2
x=-2/8
x=-1/4
KARTHIK
8 years ago
8x+4=2
8x=2-4
8x=-2
x=-2/8
x=-1/4
venky
8 years ago
X=-1/4
Gaddam Chethan
8 years ago
8x+4=2
8x=2-4
8x=-2
x=-2/8
x=-1/8
so the required answer is x=-1/4
so 8x+4=-1/4........................
Anurag anand
7 years ago
TAKING 4 TO RHS AND DIVIDING BY 8 WE GET X=
Mudit
7 years ago
8x+4=2
Shreyas Patil
7 years ago
-1/4....
Ritika Das
6 years ago
8x+4=2or, 8x=-2or,x=-2/8or,x=-1/4
Agastya Varma
6 years ago
8x+4=2
=> 8x+4-4=2-4. ( subtracting 4 on both sides)
8x=-2
=> X= -2/8.  (Dividing by 8 on both sides)
X= -1/4.
yathartha gupta
6 years ago
we have, 8x+4=2adding -4 to both side,we get(8x+4)-4=2-48x+4-4=-28x=-28x/8=-2/8x=-1/4thus the value of x=-1/4 for which the equation is satisfied.checkLHS= 8x+4 =8*(-1/4)+4 =2RHS=2Therefore left hand side=right hand
Nitesh
6 years ago
From the question ,we have 8x+4=2 =》8x=2-4=》8x=-2=》x=-2/8=》x=-1/4 Hence the answer is -1/4First try yourself by applying rule of sign change when a number or quantity is kept from one side of equality sign to another side of equality it`s sign changes from +to - .
Harshit Panwar
6 years ago
8x+4=28x=2-48x=
ankan sen
6 years ago
first see the equation(eq)
the solve like this
the question is this
8x+4=2
the first step
8x=2-4
then
8x=-2
x=-2/8
on solving this
we get
x=-1/4
Abhinav Gupta
6 years ago
we have, 8x+4=2adding -4 to both side,we get(8x+4)-4=2-48x+4-4=-28x=-28x/8=-2/8x=-1/4thus the value of x=-1/4 for which the equation is satisfied.checkLHS= 8x+4 =8*(-1/4)+4 =2RHS=2Therefore left hand side = right hand side
Vivek
6 years ago
Answer -1/48x+4=28x= -2X= -2/8X= -1/4This can also be verified by putting the value of x in the above equation.
C.Vishaal
6 years ago
8x + 4 = 2. So bring four to the other side, which gives 8x = 2 – 4 (Transposition)
That gives you , 8x = -2
                             x = -2/8
 
Which when simplifies gives you a final answer of -1/4.
 
Hope this answer helps.
Amit
6 years ago
we have, 8x+4=2adding -4 to both side,we get(8x+4)-4=2-48x+4-4=-28x=-28x/8=-2/8x=-1/4thus the value of x=-1/4 for which the equation is satisfied.checkLHS= 8x+4 =8*(-1/4)+4 =2RHS=2Therefore left hand side=right hand side 3 years ago8x+4=28x=4-2x=2/8
Amit
6 years ago
8x+4=28x=4-2x=2/8.......M...m
unnati kasaundhan
6 years ago
8x+4=28x=2-4=-2x=-2/8=-1/4So, x=-1/4
Priyansh
6 years ago
negative 0.25
sai sandeep
6 years ago
8x+4=2             this is called linear equation
8x=2-4              transforming 4 to other side
8x=-2              
x=-2/8
x=-1/4 it is negative answer which is equall to -0.25
Siddharth Singh
6 years ago
8x+4=2 ......it is a linear equation.......8x+4=28x=-2x =-1/4......Thanks...........From Siddharth Singh
Varsha Shivanand K
6 years ago
8x+4=2
8x+4-4=2-4
8x= -2
8x\8 = -2\8
x= -1/4   ( BY SYSTEMATIC METHOD)
Therefore -1/4 is the answer.
 
 
hope the answer helps you.
VARSHA SHIVANAND K.
 
SR Roy
6 years ago
8x+4=2  this is called linear equation
8x=2-4   transforming 4 to other side
8x=-2         
x=-2/8
x=-1/4 it is negative answer which is equall to -0.25
Thank you
Bharat Luthra
5 years ago
8x+4=2
or 8x+4-4=2-4
or 8x=-2
or 8x/8=-2/8
or x=-2/8
or x=-1/4=-0.25
thus, the solution of the equation is x=-0.25
Alice
5 years ago
8x+4=2
4x+2=1
4x=-1
x=-1/4
Poonam Sandhu
5 years ago
8x+4=2
8x=2-4 (transposition)
8x=-2
x=-2/8
thus x= -1/4
thus the value of x is (-1/4)
hope this answer will be helpful.
Mohd Mujtaba
5 years ago
8x+4=2.
So add -4 on both side you get ,
8x=-2
Now divide by 8 on both sides you get,
x=-1/4.
Hope it helps you.☺☺
Sai Soumya
5 years ago
=>8x+4=2
taking 4 to the other sie , we get,
=>8x=2-4
=>8x=(-2)
taking 8 to the other side , we get
=>x=(-2)/8
 
so value of x=(-1)/4
hope this help you
Nawin
5 years ago
8x+4=2
(8x+4)-4=(2)-4
8x=-2
(8x)/8=(-2)/8
x=(-1)/4(or)(-0.25)
Hope answer helped you
Bye
Cheers
Om
5 years ago
Dear student
8x+4=2.  (Given)
8x=-4                                [by taking 4 to RHS]
x=-(1/2)                             [by takin 8 to RHS]
 
I hope this answer helps you
With regards 
A fellow student
 
Colonizer
5 years ago
Answer for your question ”8x+4=2” is,
 
8x+4= 2
8x= 2-4
8x= -2
x= -2/8
x= -0.25
 
So, the answer is -0.25.
Hope you understand this answer. write if you have more questions.
Sai Soumya
5 years ago
Acoording tot the given question
 
8x+4= 2
8x= 2-4
8x= -2
x= -2/8
Hope this will help you
have a good day
yathartha gupta
5 years ago
8x+4=0
I.e. 8x= -4
I.e. x= -4÷8
I.e. x= -0.5    or         -1/2
    It is the mejor I hope you can understand it easily
 
 
john cena
5 years ago
dear student
8x+4=2
divide both side by 2
we get 
4x+2=1
subtract both side by 1
4x=-1
divide both side by 4
x=-1/4
hence the value of x=-1/4
Bharat Luthra
5 years ago
8x+4=2
=> 8x+4-4=2-4
=>8x=-2
=>8x/8=-2/8
=>x=-2/8
=>x=-1/4  Ans
See, we can add, subtract, multiply or divide with the same quantity on both sides
Sayantan Garai
5 years ago
8x+4=2
8x=2–4
8x=-2
x=–2/8
x=–1/4.
Hence the answer is x=-1/4.
It is called a linear equation in one variable.
Varsha Shivanand K
4 years ago
THE ANSWER IS AS FOLLOWS--
8x+4=2
8x=2-4
8x=-2
x=-2/8
x=-1/4
THIS IS CALLED LINEAR EQUATION IN ONE VARIABLE
 
hope it helps
Sanjai
4 years ago
8x+4=2
    8x= -2
     X = -2/8. => -1/4
  This is a simple case. Just adjust the variable and constants and make it easier for substitution
Rohit Sah
4 years ago
we have, 8x+4=2adding -4 to both side,we get(8x+4)-4=2-48x+4-4=-28x=-28x/8=-2/8x=-1/4thus the value of x=-1/4 
Swati
4 years ago
Dear Student,
 
The give equation is :- 
8x + 4 = -2
 
Now, take 4 to the RHS of the equation
8x = 2-4
implies, 8x = -2
then, x = -2/8
gives, x= -1/4 
 
Hope this will help you out.
 
With regards
Swati.